OpenStudy (anonymous):

tangent line equation of f(x)=x+x/3

OpenStudy (anonymous):

\[f(x) = x+\frac{ 1 }{ 3 }x\] Add. \[f(x) = \frac{ 4 }{ 3 }x\] Now differentiate. That's your slope. Without a specific point, it's just y=mx+b where m is the answer you got for the derivative.

OpenStudy (anonymous):

I have the specific point of (-3,-4)

OpenStudy (anonymous):

Thank you!